On Sunday 4th a problem was found with the compressors, affecting the F2 vacuum. The instrument needs to be warmed-up to inspect the issue and therefore will not be available for a week or so.
After the COVID-19 telescope shutdown, F2 will continue off the telescope until further notice, due to pending work on the instrument cold heads.
F2 will be off the telescope until March 27th, due to work on the instrument cold heads.
Update November 6th, 2019
F2 is back on sky, fully operational for imaging and long-slit spectroscopy with the OIWFS.
Update September 26th, 2019
Recovery of the cooling was unsuccessful and the cold-heads in F2 must be replaced. This means that the instrument will be unavailable for approximately 4 weeks.
Update September 12th, 2019
The F2 detector has been warming up and has reached the limit for operation. We are working to recover the cooling, but until this is fixed F2 is not available for observations.